NASA's Parker Solar Probe is currently en route to the outer corona of the sun. At its closest approach it will be traveling at 7.00 x 105 km/hr. How fast will it be going in units of mi/s? 5,280 feet = 1 mile
Answer #1

1hr = 3600 s

Thus, 7.00×105 Km/hr = 7.00×105/3600 Km/s

= 1.94×102 Km/s

Now, 1mi = 1.609 Km

So, 1.94×102 Km/s = 1.94×102/1.609 mi/s

= 120.8 mi/s